# A generalization of the relaxation-based matrix splitting iterative method for solving the system of generalized absolute value equations

### v1

@@ 2 The GRMS iterative method
    If $M=\dfrac{1}{2}\left(\alpha_{\rm FPI\text{-}SS}I+A\right)$ with $\alpha_{\rm FPI\text{-}SS}>0$, $Q=I,Q_{1}=\frac{\theta}{\omega_{\rm FPI\text{-}SS}}I,~{}H=0$, then the GRMS iteration (2.3) turns into the shift-splitting fixed point iteration (FPI-SS) [15]
    $\begin{cases}x^{(k+1)}=(\alpha_{\rm FPI\text{-}SS}I+A)^{-1}(\alpha_{\rm FPI\text{-}SS}I-A)x^{(k)}+2(\alpha_{\rm FPI\text{-}SS}I+A)^{-1}(By^{(k)}+b),\\ y^{(k+1)}=(1-\omega_{\rm FPI\text{-}SS})y^{(k)}+\omega_{\rm FPI\text{-}SS}|x^{(k+1)}|,\end{cases}$ (2.7)
    where $\alpha_{\rm FPI\text{-}SS}$ is a positive constant such that $\alpha_{\rm FPI\text{-}SS}I+A$ is nonsingular and $\omega_{\rm FPI\text{-}SS}$ is a positive constant. Particularly, when $M=A$, $Q=I,Q_{1}=\frac{\theta}{\omega_{\rm FPI}}I$ and $H=0$, the GRMS iteration (2.3) turns into the fixed point iteration (FPI) for solving the GAVEs (1.1) [15]
    $\begin{cases}x^{(k+1)}=A^{-1}(By^{(k)}+b),\\ y^{(k+1)}=(1-\omega_{\rm FPI})y^{(k)}+\omega_{\rm FPI}|x^{(k+1)}|,\end{cases}$ (2.8)
    where $\omega_{\rm FPI}$ is a positive constant. Indeed, (2.8) is a generalization of (1.8).
@@
    If $M=\frac{1}{\alpha_{\rm MGSOR}}A$, $Q=Q_{\rm MGSOR},Q_{1}=\frac{\theta}{\beta_{\rm MGSOR}}Q_{\rm MGSOR}$ and $H=0$, then the GRMS iteration (2.3) becomes the MGSOR iteration
    $\begin{cases}x^{(k+1)}=(1-\alpha_{\rm MGSOR})x^{(k)}+\alpha_{\rm MGSOR}A^{-1}\left(BQ_{\rm MGSOR}y^{(k)}+b\right),\\ y^{(k+1)}=(1-\beta_{\rm MGSOR})y^{(k)}+\beta_{\rm MGSOR}Q^{-1}_{\rm MGSOR}|x^{(k+1)}|\end{cases}$ (2.9)
    for solving the GAVEs (1.1) with $\alpha_{\rm MGSOR}>0$ and $\beta_{\rm MGSOR}>0$, which is an extension of (2.5). Specially, similar to the case for the AVEs (1.2), if $\alpha_{\rm MGSOR}=\beta_{\rm MGSOR}=\omega_{\rm SOR}>0$ and $Q_{\rm MGSOR}=I$, the MGSOR iteration (2.9) becomes the SOR-like iteration
    $\begin{cases}x^{(k+1)}=(1-\omega_{\rm SOR})x^{(k)}+\omega_{\rm SOR}A^{-1}\left(By^{(k)}+b\right),\\ y^{(k+1)}=(1-\omega_{\rm SOR})y^{(k)}+\omega_{\rm SOR}|x^{(k+1)}|.\end{cases}$ (2.10)
    for solving the GAVEs (1.1), which is a generalization of (1.7). If $\alpha_{\rm MGSOR}=1$, $\beta_{\rm MGSOR}=\tau_{\rm MFPI}>0$ and $Q_{\rm MGSOR}=Q_{\rm MFPI}$, the MGSOR iteration (2.9) turns into the following MFPI
    $\begin{cases}x^{(k+1)}=A^{-1}(BQ_{\rm MFPI}y^{(k)}+b),\\ y^{(k+1)}=(1-\tau_{\rm MFPI})y^{(k)}+\tau_{\rm MFPI}Q^{-1}_{\rm MFPI}|x^{(k+1)}|\end{cases}$ (2.11)
    for solving the GAVEs (1.1), which is a generalization of (1.10).
@@
    Let $B\in\mathbb{R}^{n\times n}$ be nonsingular. If $M=\frac{1}{\alpha_{\rm MSOR}\omega_{\rm MSOR}}A$, $Q=I$, $Q_{1}=\frac{\theta(2-\omega_{\rm MSOR})}{2\omega_{\rm MSOR}}B^{-1}$ and $H=\frac{\theta(1-\alpha_{\rm MSOR}\omega_{\rm MSOR})}{\alpha_{\rm MSOR}\omega_{\rm MSOR}}B^{-1}A$ ($\alpha_{\rm MSOR},\omega_{\rm MSOR}>0$ and $\omega_{\rm MSOR}\neq 2$), the GRMS iteration (2.3) becomes the MSOR iteration
    $\begin{cases}x^{(k+1)}=(1-\alpha_{\rm MSOR}\omega_{\rm MSOR})x^{(k)}+\alpha_{\rm MSOR}\omega_{\rm MSOR}A^{-1}(By^{(k)}+b),\\ y^{(k+1)}=y^{(k)}+\frac{2\omega_{\rm MSOR}}{2-\omega_{\rm MSOR}}(-Ax^{(k+1)}+B|x^{(k+1)}|+b)\end{cases}$ (2.12)
    for solving the GAVEs (1.1), which is a generalization of (2.4). Particularly, if $\alpha_{\rm MSOR}=\frac{2\sigma_{\rm NSOR}+\omega_{\rm NSOR}}{2}$ and $\omega_{\rm MSOR}=\frac{2\omega_{\rm NSOR}}{2\sigma_{\rm NSOR}+\omega_{\rm NSOR}}$, the MSOR iteration (2.12) reduces to the NSOR iteration
    $\begin{cases}x^{(k+1)}=x^{(k)}+\omega_{\rm NSOR}A^{-1}(b-Ax^{(k)}+By^{(k)}),\\ y^{(k+1)}=y^{(k)}+\dfrac{\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}(b+B|x^{(k+1)}|-Ax^{(k+1)})\end{cases}$ (2.13)
    for solving the GAVEs (1.1) with $\omega_{\rm NSOR}>0$ and $\sigma_{\rm NSOR}>0$, which is a generalization of (1.9).
@@
    Let $B$ be nonsingular. If $\Omega_{\rm MAMS}$ is a positive diagonal matrix such that $M=M_{\rm MAMS}+\Omega_{\rm MAMS}$ is nonsingular, where $A=M_{\rm MAMS}-N_{\rm MAMS}$. Furthermore, let $Q=I$, $Q_{1}=\theta I$, and $H=\theta\beta_{\rm MAMS}B^{-1}$ with $\beta_{\rm MAMS}\in\mathbb{R}$. Then the GRMS iteration (2.3) turns into the following MAMS iteration [32]
    $x^{(k+2)}=(M_{\rm MAMS}+\Omega_{\rm MAMS})^{-1}\left[\big{(}N_{\rm MAMS}+\Omega_{\rm MAMS}\big{)}x^{(k+1)}+B|x^{(k+1)}|+\beta_{\rm MAMS}\big{(}x^{(k+1)}-x^{(k)}\big{)}+b\right]$ (2.14)
@@ 3 Convergence analysis
In this section, we will analyse the convergence of the GRMS iteration (2.3). Indeed, we have the following convergence theorem.
Theorem 3.1.
Assume that $A,B,H\in\mathbb{R}^{n\times n}$ and $Q\in\mathbb{R}^{n\times n}$ is nonsingular. Let $A=M-N$ and $Q=Q_{1}-Q_{2}$ with $M$ and $Q_{1}$ being nonsingular. Denote
$a=\|M^{-1}N\|,c=\|M^{-1}BQ\|,d=\|Q_{1}^{-1}Q_{2}\|,\alpha=\|Q_{1}^{-1}\|,\beta=\|Q_{1}^{-1}H\|.$ (3.1)
@@
$\big{|}a|1-\theta|+\theta ad-c\beta\big{|}<1$ (3.2)
@@
$c(\theta\alpha+2\beta)<(1-a)(1-|1-\theta|-\theta d),$ (3.3)
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the GRMS iteration (2.3)converges to $(x^{*},y^{*}=Q^{-1}|x^{*}|)$.
@@
then from (2.3), for $k\geq 1$, we have
$\begin{cases}e^{(k+1)}_{x}=M^{-1}Ne^{(k)}_{x}+M^{-1}BQe^{(k)}_{y},\\ e^{(k+1)}_{y}=(1-\theta)e^{(k)}_{y}+\theta Q_{1}^{-1}Q_{2}e^{(k)}_{y}+\theta Q_{1}^{-1}(|x^{(k+1)}|-|x^{(k)}|)+Q_{1}^{-1}He^{(k+1)}_{x}-Q_{1}^{-1}He^{(k)}_{x}.\end{cases}$ (3.4)
It follows from (3.1) and (3.4) that
$\begin{cases}\|e^{(k+1)}_{x}\|\leq a\|e^{(k)}_{x}\|+c\|e^{(k)}_{y}\|,\\ \|e^{(k+1)}_{y}\|\leq\big{(}|1-\theta|+\theta d\big{)}\|e^{(k)}_{y}\|+\big{(}\theta\alpha+\beta\big{)}\|e^{(k+1)}_{x}\|+\beta\|e^{(k)}_{x}\|,\end{cases}$ (3.5)
in which the inequality $\||x^{(k+1)}|-|x^{(k)}|\|\leq\|x^{(k+1)}-x^{(k)}\|$ is used. It follows from (3.5) that
$\begin{bmatrix}1&0\\ -(\theta\alpha+\beta)&1\end{bmatrix}\begin{bmatrix}\|e_{x}^{(k+1)}\|\\ \|e_{y}^{(k+1)}\|\end{bmatrix}\leq\begin{bmatrix}a&c\\ \beta&|1-\theta|+\theta d\end{bmatrix}\begin{bmatrix}\|e_{x}^{(k)}\|\\ \|e_{y}^{(k)}\|\end{bmatrix}.$ (3.6)
Left multiplying both sides of the inequality (3.6) by the nonnegative matrix $\begin{bmatrix}1&0\\ \theta\alpha+\beta&1\end{bmatrix}$, we have
$\begin{bmatrix}\|e_{x}^{(k+1)}\|\\ \|e_{y}^{(k+1)}\|\end{bmatrix}\leq T\begin{bmatrix}\|e_{x}^{(k)}\|\\ \|e_{y}^{(k)}\|\end{bmatrix},$ (3.7)
@@
$T=\begin{bmatrix}a&c\\ a(\theta\alpha+\beta)+\beta&c(\theta\alpha+\beta)+\theta d+|1-\theta|\end{bmatrix}.$ (3.8)
From (3.7), we can deduce that
$\begin{bmatrix}\|e_{x}^{(k+1)}\|\\ \|e_{y}^{(k+1)}\|\end{bmatrix}\leq T\begin{bmatrix}\|e_{x}^{(k)}\|\\ \|e_{y}^{(k)}\|\end{bmatrix}\leq\dots\leq T^{k}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}.$ (3.9)
Then for $k\geq 1$ and for any $m\in\mathbb{N^{+}}$, if $\rho(T)<1$, it follows from (3.9) that
$\begin{bmatrix}\|x^{(k+m)}-x^{(k)}\|\\ \|y^{(k+m)}-y^{(k)}\|\end{bmatrix}$ $=\begin{bmatrix}\Big{\|}\sum\limits_{j=0}^{m-1}e^{(k+j+1)}_{x}\Big{\|}\\ \Big{\|}\sum\limits_{j=0}^{m-1}e_{y}^{(k+j+1)}\Big{\|}\end{bmatrix}\leq\begin{bmatrix}\sum\limits_{j=0}^{m-1}\|e^{(k+j+1)}_{x}\|\\ \sum\limits_{j=0}^{m-1}\|e_{y}^{(k+j+1)}\|\end{bmatrix}=\sum\limits_{j=0}^{m-1}\begin{bmatrix}\|e^{(k+j+1)}_{x}\|\\ \|e_{y}^{(k+j+1)}\|\end{bmatrix}$ (3.18)
$\leq\left(\sum_{j=0}^{m-1}T^{j}\right)\left(\begin{bmatrix}\|e^{(k+1)}_{x}\|\\ \|e_{y}^{(k+1)}\|\end{bmatrix}\right)\leq\left(\sum_{j=0}^{m-1}T^{j}\right)\left(T^{k}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}\right)$ (3.23)
$\leq\left(\sum_{j=0}^{\infty}T^{j}\right)\left(T^{k}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}\right)=(I-T)^{-1}T^{k}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}.$ (3.28)
In addition, if $\rho(T)<1$, then $\lim\limits_{k\rightarrow\infty}T^{k}=0$. Hence, for $\forall\varepsilon>0$, there exist $K\in\mathbb{N}^{+}$ such that for $k>K$ we obtain
$|T^{k}|<\begin{bmatrix}\varepsilon&\varepsilon\\ \varepsilon&\varepsilon\end{bmatrix}.$ (3.29)
Moreover, for $(I-T)^{-1}$ and $\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}$, there is an $M>0$ such that
$|(I-T)^{-1}|<\begin{bmatrix}M&M\\ M&M\end{bmatrix}\quad\text{and}\quad\bigg{|}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}\bigg{|}<\begin{bmatrix}M\\ M\end{bmatrix}.$ (3.30)
According to (3.28), (3.29) and (3.30), if $\rho(T)<1$, then for all $\hat{\varepsilon}=4M^{2}\varepsilon>0$, $\exists K>0$ (as mentioned above) such that for all $k>K$ and any $m\in\mathbb{N}^{+}$ we have
$\begin{bmatrix}\|x^{(k+m)}-x^{(k)}\|\\ \|y^{(k+m)}-y^{(k)}\|\end{bmatrix}=\left|\begin{bmatrix}\|x^{(k+m)}-x^{(k)}\|\\ \|y^{(k+m)}-y^{(k)}\|\end{bmatrix}\right|\leq\left|(I-T)^{-1}\right|\left|T^{k}\right|\left|\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}\right|<\begin{bmatrix}\hat{\varepsilon}\\ \hat{\varepsilon}\\ \end{bmatrix},$ (3.31)
from which we can conclude that both $\{x^{(k)}\}^{\infty}_{k=0}$ and $\{y^{(k)}\}^{\infty}_{k=0}$ are Cauchy sequences. Since the Cauchy sequences $\{x^{(k)}\}^{\infty}_{k=0}\subseteq\mathbb{R}^{n}$ and $\{y^{(k)}\}^{\infty}_{k=0}\subseteq\mathbb{R}^{n}$, they are convergent. Hence, we can let $\mathop{\lim}\limits_{k\rightarrow\infty}x^{(k)}=x^{*}$ and $\mathop{\lim}\limits_{k\rightarrow\infty}y^{(k)}=y^{*}$. Then it follows from (2.3) that
@@
$\begin{cases}Ax^{*}-B|x^{*}|-b=0,\\ y^{*}=Q^{-1}|x^{*}|.\end{cases}$ (3.32)
The first equation of (3.32) means that $x^{*}$ is a solution of the GAVEs (1.1).
In the following, we will prove that $\rho(T)<1$ if (3.2) and (3.3) hold. Suppose that $\lambda$ is an eigenvalue of $T$, from (3.8), we have
@@
Using [28, Lemma 2.1], $\rho(T)<1$ provided that
@@
In other words, (3.2) and (3.3) implies $\rho(T)<1$.
The final step is to prove the uniqueness of the solution to the GAVEs (1.1). Assume that there is another solution $\bar{x}^{*}$ to the GAVEs (1.1), then we can uniquely determine $\bar{y}^{*}=Q^{-1}|\bar{x}^{*}|$. Using (2.2), we have
$x^{*}-\bar{x}^{*}$ $=M^{-1}N(x^{*}-\bar{x}^{*})+M^{-1}BQ(y^{*}-\bar{y}^{*}),$ (3.33)
@@
It follows from (3.33) and (3.1) that
$\|x^{*}-\bar{x}^{*}\|\leq a\|x^{*}-\bar{x}^{*}\|+c\|y^{*}-\bar{y}^{*}\|,$ (3.34)
$\|y^{*}-\bar{y}^{*}\|\leq(|1-\theta|+\theta d)\|y^{*}-\bar{y}^{*}\|+(\theta\alpha+2\beta)\|x^{*}-\bar{x}^{*}\|.$ (3.35)
If (3.2) and (3.3) hold, we can conclude that $a<1~{}\text{and}~{}|1-\theta|+\theta d<1$. In fact, the condition (3.3) implies that either

### v2

@@ 2 The GRMS iterative method
    If $M=\dfrac{1}{2}\left(\alpha_{\rm FPI\text{-}SS}I+A\right)$ with $\alpha_{\rm FPI\text{-}SS}>0$, $Q=I,Q_{1}=\frac{\theta}{\omega_{\rm FPI\text{-}SS}}I,~{}H=0$, then the GRMS iteration (13) turns into the shift-splitting fixed point iteration (FPI-SS) [16]
    $\begin{cases}x^{(k+1)}=(\alpha_{\rm FPI\text{-}SS}I+A)^{-1}(\alpha_{\rm FPI\text{-}SS}I-A)x^{(k)}+2(\alpha_{\rm FPI\text{-}SS}I+A)^{-1}(By^{(k)}+b),\\ y^{(k+1)}=(1-\omega_{\rm FPI\text{-}SS})y^{(k)}+\omega_{\rm FPI\text{-}SS}|x^{(k+1)}|,\end{cases}$ (17)
    where $\alpha_{\rm FPI\text{-}SS}$ is a positive constant such that $\alpha_{\rm FPI\text{-}SS}I+A$ is nonsingular and $\omega_{\rm FPI\text{-}SS}$ is a positive constant. Particularly, when $M=A$, $Q=I,Q_{1}=\frac{\theta}{\omega_{\rm FPI}}I$ and $H=0$, the GRMS iteration (13) turns into the fixed point iteration (FPI) for solving GAVEs (1) [16]
    $\begin{cases}x^{(k+1)}=A^{-1}(By^{(k)}+b),\\ y^{(k+1)}=(1-\omega_{\rm FPI})y^{(k)}+\omega_{\rm FPI}|x^{(k+1)}|,\end{cases}$ (18)
    where $\omega_{\rm FPI}$ is a positive constant. Indeed, (18) is a generalization of (8).
@@
    If $M=\frac{1}{\alpha_{\rm MGSOR}}A$, $Q=Q_{\rm MGSOR},Q_{1}=\frac{\theta}{\beta_{\rm MGSOR}}Q_{\rm MGSOR}$ and $H=0$, then the GRMS iteration (13) becomes the MGSOR iteration
    $\begin{cases}x^{(k+1)}=(1-\alpha_{\rm MGSOR})x^{(k)}+\alpha_{\rm MGSOR}A^{-1}\left(BQ_{\rm MGSOR}y^{(k)}+b\right),\\ y^{(k+1)}=(1-\beta_{\rm MGSOR})y^{(k)}+\beta_{\rm MGSOR}Q^{-1}_{\rm MGSOR}|x^{(k+1)}|\end{cases}$ (19)
    for solving GAVEs (1) with $\alpha_{\rm MGSOR}>0$ and $\beta_{\rm MGSOR}>0$, which is an extension of (15). Specially, similar to the case for AVEs (2), if $\alpha_{\rm MGSOR}=\beta_{\rm MGSOR}=\omega_{\rm SOR}>0$ and $Q_{\rm MGSOR}=I$, the MGSOR iteration (19) becomes the SOR-like iteration
    $\begin{cases}x^{(k+1)}=(1-\omega_{\rm SOR})x^{(k)}+\omega_{\rm SOR}A^{-1}\left(By^{(k)}+b\right),\\ y^{(k+1)}=(1-\omega_{\rm SOR})y^{(k)}+\omega_{\rm SOR}|x^{(k+1)}|\end{cases}$ (20)
    for solving GAVEs (1), which is a generalization of (7). If $\alpha_{\rm MGSOR}=1$, $\beta_{\rm MGSOR}=\tau_{\rm MFPI}>0$ and $Q_{\rm MGSOR}=Q_{\rm MFPI}$, the MGSOR iteration (19) turns into the following MFPI
    $\begin{cases}x^{(k+1)}=A^{-1}(BQ_{\rm MFPI}y^{(k)}+b),\\ y^{(k+1)}=(1-\tau_{\rm MFPI})y^{(k)}+\tau_{\rm MFPI}Q^{-1}_{\rm MFPI}|x^{(k+1)}|\end{cases}$ (21)
    for solving GAVEs (1), which is a generalization of (10).
@@
    Let $B$ be nonsingular. If $M=\frac{1}{\alpha_{\rm MSOR}\omega_{\rm MSOR}}A$, $Q=I$, $Q_{1}=\frac{\theta(2-\omega_{\rm MSOR})}{2\omega_{\rm MSOR}}B^{-1}$ and $H=\frac{\theta(1-\alpha_{\rm MSOR}\omega_{\rm MSOR})}{\alpha_{\rm MSOR}\omega_{\rm MSOR}}B^{-1}A$ ($\alpha_{\rm MSOR},\omega_{\rm MSOR}>0$ and $\omega_{\rm MSOR}\neq 2$), the GRMS iteration (13) becomes the MSOR iteration
    $\begin{cases}x^{(k+1)}=(1-\alpha_{\rm MSOR}\omega_{\rm MSOR})x^{(k)}+\alpha_{\rm MSOR}\omega_{\rm MSOR}A^{-1}(By^{(k)}+b),\\ y^{(k+1)}=y^{(k)}+\frac{2\omega_{\rm MSOR}}{2-\omega_{\rm MSOR}}(-Ax^{(k+1)}+B|x^{(k+1)}|+b)\end{cases}$ (22)
    for solving GAVEs (1), which is a generalization of (14). Particularly, if $\alpha_{\rm MSOR}=\frac{2\sigma_{\rm NSOR}+\omega_{\rm NSOR}}{2}$ and $\omega_{\rm MSOR}=\frac{2\omega_{\rm NSOR}}{2\sigma_{\rm NSOR}+\omega_{\rm NSOR}}$, the MSOR iteration (22) reduces to the NSOR iteration
    $\begin{cases}x^{(k+1)}=x^{(k)}+\omega_{\rm NSOR}A^{-1}(b-Ax^{(k)}+By^{(k)}),\\ y^{(k+1)}=y^{(k)}+\dfrac{\omega_{\rm NSOR}}{\sigma_{\rm NSOR}}(b+B|x^{(k+1)}|-Ax^{(k+1)})\end{cases}$ (23)
    for solving GAVEs (1) with $\omega_{\rm NSOR}>0$ and $\sigma_{\rm NSOR}>0$, which is a generalization of (9).
@@
    Let $B$ be nonsingular. If $\Omega_{\rm MAMS}$ is a positive diagonal matrix such that $M=M_{\rm MAMS}+\Omega_{\rm MAMS}$ is nonsingular, where $A=M_{\rm MAMS}-N_{\rm MAMS}$. Furthermore, let $Q=I$, $Q_{1}=\theta I$, and $H=\theta\beta_{\rm MAMS}B^{-1}$ with $\beta_{\rm MAMS}$. Then the GRMS iteration (13) turns into the following MAMS iteration [33]
    $x^{(k+2)}=(M_{\rm MAMS}+\Omega_{\rm MAMS})^{-1}\left[\big{(}N_{\rm MAMS}+\Omega_{\rm MAMS}\big{)}x^{(k+1)}+B|x^{(k+1)}|+\beta_{\rm MAMS}\big{(}x^{(k+1)}-x^{(k)}\big{)}+b\right]$ (24)
@@ 3 Convergence analysis
In this section, we will analyse the convergence of the GRMS iteration (13). Indeed, we have the following convergence theorem.
Theorem 1.
Assume that $A,B,H\in\mathbb{R}^{n\times n}$, and $Q\in\mathbb{R}^{n\times n}$ is nonsingular. Let $A=M-N$ and $Q=Q_{1}-Q_{2}$ with $M$ and $Q_{1}$ being nonsingular. Denote
$a=\|M^{-1}N\|,c=\|M^{-1}BQ\|,d=\|Q_{1}^{-1}Q_{2}\|,\alpha=\|Q_{1}^{-1}\|,\beta=\|Q_{1}^{-1}H\|.$ (25)
@@
$\big{|}a|1-\theta|+\theta ad-c\beta\big{|}<1$ (26)
@@
$c(\theta\alpha+2\beta)<(1-a)(1-|1-\theta|-\theta d),$ (27)
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the GRMS iteration (13)converges to $(x^{*},y^{*}=Q^{-1}|x^{*}|)$.
@@
and then from (13), for $k\geq 1$, we have
$\begin{cases}e^{(k+1)}_{x}=M^{-1}Ne^{(k)}_{x}+M^{-1}BQe^{(k)}_{y},\\ e^{(k+1)}_{y}=(1-\theta)e^{(k)}_{y}+\theta Q_{1}^{-1}Q_{2}e^{(k)}_{y}+\theta Q_{1}^{-1}(|x^{(k+1)}|-|x^{(k)}|)+Q_{1}^{-1}He^{(k+1)}_{x}-Q_{1}^{-1}He^{(k)}_{x}.\end{cases}$ (28)
It follows from (25) and (28) that
$\begin{cases}\|e^{(k+1)}_{x}\|\leq a\|e^{(k)}_{x}\|+c\|e^{(k)}_{y}\|,\\ \|e^{(k+1)}_{y}\|\leq\big{(}|1-\theta|+\theta d\big{)}\|e^{(k)}_{y}\|+\big{(}\theta\alpha+\beta\big{)}\|e^{(k+1)}_{x}\|+\beta\|e^{(k)}_{x}\|,\end{cases}$ (29)
in which the inequality $\||x^{(k+1)}|-|x^{(k)}|\|\leq\|x^{(k+1)}-x^{(k)}\|$ is used. It follows from (29) that
$\begin{bmatrix}1&0\\ -(\theta\alpha+\beta)&1\end{bmatrix}\begin{bmatrix}\|e_{x}^{(k+1)}\|\\ \|e_{y}^{(k+1)}\|\end{bmatrix}\leq\begin{bmatrix}a&c\\ \beta&|1-\theta|+\theta d\end{bmatrix}\begin{bmatrix}\|e_{x}^{(k)}\|\\ \|e_{y}^{(k)}\|\end{bmatrix}.$ (30)
Left multiplying both sides of the inequality (30) by the nonnegative matrix $\begin{bmatrix}1&0\\ \theta\alpha+\beta&1\end{bmatrix}$, we have
$\begin{bmatrix}\|e_{x}^{(k+1)}\|\\ \|e_{y}^{(k+1)}\|\end{bmatrix}\leq T\begin{bmatrix}\|e_{x}^{(k)}\|\\ \|e_{y}^{(k)}\|\end{bmatrix},$ (31)
@@
$T=\begin{bmatrix}a&c\\ a(\theta\alpha+\beta)+\beta&c(\theta\alpha+\beta)+\theta d+|1-\theta|\end{bmatrix}.$ (32)
From (31), we can deduce that
$\begin{bmatrix}\|e_{x}^{(k+1)}\|\\ \|e_{y}^{(k+1)}\|\end{bmatrix}\leq T\begin{bmatrix}\|e_{x}^{(k)}\|\\ \|e_{y}^{(k)}\|\end{bmatrix}\leq\dots\leq T^{k}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}.$ (33)
Then for $k\geq 1$ and for any $m\in\mathbb{N^{+}}$, if $\rho(T)<1$, it follows from (33) that
$\begin{bmatrix}\|x^{(k+m)}-x^{(k)}\|\\ \|y^{(k+m)}-y^{(k)}\|\end{bmatrix}$ $=\begin{bmatrix}\Big{\|}\sum\limits_{j=0}^{m-1}e^{(k+j+1)}_{x}\Big{\|}\\ \Big{\|}\sum\limits_{j=0}^{m-1}e_{y}^{(k+j+1)}\Big{\|}\end{bmatrix}\leq\begin{bmatrix}\sum\limits_{j=0}^{m-1}\|e^{(k+j+1)}_{x}\|\\ \sum\limits_{j=0}^{m-1}\|e_{y}^{(k+j+1)}\|\end{bmatrix}=\sum\limits_{j=0}^{m-1}\begin{bmatrix}\|e^{(k+j+1)}_{x}\|\\ \|e_{y}^{(k+j+1)}\|\end{bmatrix}$
$\leq\left(\sum_{j=0}^{m-1}T^{j}\right)\left(\begin{bmatrix}\|e^{(k+1)}_{x}\|\\ \|e_{y}^{(k+1)}\|\end{bmatrix}\right)\leq\left(\sum_{j=0}^{m-1}T^{j}\right)\left(T^{k}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}\right)$
$\leq\left(\sum_{j=0}^{\infty}T^{j}\right)\left(T^{k}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}\right)=(I-T)^{-1}T^{k}\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}.$ (34)
In addition, if $\rho(T)<1$, and then $\lim\limits_{k\rightarrow\infty}T^{k}=0$. Hence, for $\forall\varepsilon>0$, there exist $K\in\mathbb{N}^{+}$ such that for $k>K$ we obtain
$|T^{k}|<\begin{bmatrix}\varepsilon&\varepsilon\\ \varepsilon&\varepsilon\end{bmatrix}.$ (35)
Moreover, for $(I-T)^{-1}$ and $\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}$, there is an $s>0$ such that
$|(I-T)^{-1}|<\begin{bmatrix}s&s\\ s&s\end{bmatrix}\quad\text{and}\quad\left|\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}\right|<\begin{bmatrix}s\\ s\end{bmatrix}.$ (36)
According to (34), (35) and (36), if $\rho(T)<1$, and then for all $\hat{\varepsilon}=4s^{2}\varepsilon>0$, $\exists K>0$ (as mentioned above) such that for all $k>K$ and any $m\in\mathbb{N}^{+}$ we have
$\begin{bmatrix}\|x^{(k+m)}-x^{(k)}\|\\ \|y^{(k+m)}-y^{(k)}\|\end{bmatrix}=\left|\begin{bmatrix}\|x^{(k+m)}-x^{(k)}\|\\ \|y^{(k+m)}-y^{(k)}\|\end{bmatrix}\right|\leq\left|(I-T)^{-1}\right|\left|T^{k}\right|\left|\begin{bmatrix}\|e_{x}^{(1)}\|\\ \|e_{y}^{(1)}\|\end{bmatrix}\right|<\begin{bmatrix}\hat{\varepsilon}\\ \hat{\varepsilon}\\ \end{bmatrix},$ (37)
from which we can conclude that both $\{x^{(k)}\}^{\infty}_{k=0}$ and $\{y^{(k)}\}^{\infty}_{k=0}$ are Cauchy sequences. Since the Cauchy sequences $\{x^{(k)}\}^{\infty}_{k=0}\subseteq\mathbb{R}^{n}$ and $\{y^{(k)}\}^{\infty}_{k=0}\subseteq\mathbb{R}^{n}$, they are convergent. Hence, we can let $\mathop{\lim}\limits_{k\rightarrow\infty}x^{(k)}=x^{*}$ and $\mathop{\lim}\limits_{k\rightarrow\infty}y^{(k)}=y^{*}$. Then it follows from (13) that
@@
$\begin{cases}Ax^{*}-B|x^{*}|-b=0,\\ y^{*}=Q^{-1}|x^{*}|.\end{cases}$ (38)
The first equation of (38) means that $x^{*}$ is a solution of GAVEs (1).
In the following, we will prove that $\rho(T)<1$ if (26) and (27) hold. Suppose that $\lambda$ is an eigenvalue of $T$, from (32), we have
@@
Using [29, Lemma 2.1], $\rho(T)<1$ provided that
@@
In other words, (26) and (27) implies $\rho(T)<1$.
The final step is to prove the uniqueness of the solution to GAVEs (1). Assume that there is another solution $\bar{x}^{*}$ to GAVEs (1), and then we can uniquely determine $\bar{y}^{*}=Q^{-1}|\bar{x}^{*}|$. Using (12), we have
$x^{*}-\bar{x}^{*}$ $=M^{-1}N(x^{*}-\bar{x}^{*})+M^{-1}BQ(y^{*}-\bar{y}^{*}),$ (39)
@@
It follows from (39) and (25) that
$\|x^{*}-\bar{x}^{*}\|\leq a\|x^{*}-\bar{x}^{*}\|+c\|y^{*}-\bar{y}^{*}\|,$ (40)
$\|y^{*}-\bar{y}^{*}\|\leq(|1-\theta|+\theta d)\|y^{*}-\bar{y}^{*}\|+(\theta\alpha+2\beta)\|x^{*}-\bar{x}^{*}\|.$ (41)
If (26) and (27) hold, we can conclude that $a<1~{}\text{and}~{}|1-\theta|+\theta d<1$. In fact, the condition (27) implies that either